Hickory Company manufactures two products—14,000 units of Product Y and 6,000 units of Product Z. The company uses a plantwide overhead rate based on direct labor-hours. It is considering implementing an activity-based costing (ABC) system that allocates all of its manufacturing overhead to four cost pools. The following additional information is available for the company as a whole and for Products Y and Z: Activity Cost Pool Activity Measure Estimated Overhead Cost Expected Activity Machining Machine-hours $ 198,000 10,000 MHs Machine setups Number of setups $ 86,400 180 setups Production design Number of products $ 82,000 2 products General factory Direct labor-hours $ 248,000 12,000 DLHs Activity Measure Product Y Product Z Machine-hours 6,800 3,200 Number of setups 50 130 Number of products 1 1 Direct labor-hours 7,800 4,200

11. Using the plantwide overhead rate, what percentage of the total overhead cost is allocated to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your "Percentage" answers to 2 decimal places. (i.e. 0.1234 should be entered as 12.34).)

12.

Using the ABC system, what percentage of the Machining costs is assigned to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your "Percentage" answers to 2 decimal places. (i.e. 0.1234 should be entered as 12.34).)

13. Using the ABC system, what percentage of Machine Setups cost is assigned to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your "Percentage" answers to 2 decimal places. (i.e. 0.1234 should be entered as 12.34).)

14. Using the ABC system, what percentage of the Product Design cost is assigned to Product Y and Product Z? (Round your "Percentage" answers to 2 decimal places. (i.e. 0.1234 should be entered as 12.34)).

15. Using the ABC system, what percentage of the General Factory cost is assigned to Product Y and Product Z? (Round your intermediate calculations to 2 decimal places. Round your "Percentage" answers to 2 decimal places. (i.e. 0.1234 should be entered as 12.34).)
Answer #1
11
Percentage of the total overhead cost:
Product Y 65.00% =7800/12000
Product Z 35.00% =4200/12000
12
Percentage of the Machining costs:
Product Y 68.00% =6800/10000
Product Z 32.00% =3200/10000
13
Percentage of Machine Setups cost:
Product Y 27.78% =50/180
Product Z 72.22% =130/180
14
Percentage of the Product Design cost:
Product Y 50.00%
Product Z 50.00%
15
Percentage of the General Factory cost:
Product Y 65.00% =7800/12000
Product Z 35.00% =4200/12000